Solve for Y(s), the Laplace transform of the solution y(t) to the initial value problem below y-7y +10y=te2 y(0)= 3. y(0)-
Answer #1

| yl_7y + loy - test 710): 3, ylo)--3 We have the following form formulas, 2 [f (+] - 5 L{ fH)} - 5 1760) – sn-2a1(6) - 5Fol yl - y has an obler n=1 {y} sasflt) -st1710) LE4B - SF(s) -3 7 6 For y. so has an older n=0) Ego] SO LEF(+23. 6 (L84 - 1Lşte2th - tead 2 2t We have all the Laplane the transtohus and let us plug in the equato tylt 10y- tett {{y_72{y} + 10thy LF(s) = 1 + 36 - 24 5-23- 36-5X5-2 535-3 Same 1 st 7 (s) are 715-20 to 24 at 35 (3-532